OpenStudy (anonymous):

Value of cos^2 (48) - sin^2 (12)?

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (anonymous):

@.Sam. @Callisto @jhonyy9 @myko @Mani_Jha @experimentX @eliassaab @inkyvoyd @maheshmeghwal9

mathslover (mathslover):

i am solving it further wait

OpenStudy (anonymous):

ok

OpenStudy (maheshmeghwal9):

This question is using formula: - \[\cos 2\theta and \cos2(2\theta).\]

OpenStudy (anonymous):

cos^2=1-sin^2 and sin^2=1-cos^2 so cos^2(48)=1-sin^2(48) and sin^2=1-cos^2(12) 1-sin^2(48)-(1-cos^2(12)) =1-sin^2(48)-1-cos^2(12) =-sin^2(48)-cos^2(12)...

OpenStudy (maheshmeghwal9):

@mathslover will definitely do it:)

OpenStudy (maheshmeghwal9):

if u take theta= 12

OpenStudy (anonymous):

the answer should be